OpenStudy (anonymous):

The cross-section of a tunnel has the form of a rectangle surmounted by a semi-circle. The perimeter of this cross section is 18 meters. For what radius of the semi-circle will the cross-section have maximum area?

OpenStudy (anonymous):

you need to take 18 the perimeter including the semi-circle or just the perimeter of the rectangle?

OpenStudy (anonymous):

edit: is 18 the perimiter of the whole thing semi-circle included or the perimeter of the box?

OpenStudy (anonymous):

its the perimeter of the cross section, but i don't really know how the image looks

OpenStudy (zehanz):

This is what it looks:

OpenStudy (zehanz):

Now you have two variables, r and h (see image). Because we know the perimeter, we can eliminate h. The perimeter is 2r + 2h + halfcircle. Because the perimeter of a circle is 2pi*r, a halfcircle has perimeter pi*r. So: 18 = 2r + 2h + pi*r, so 2h = 18 - 2r - pi*r, and \(h=9-r -\frac{\pi}{2}r=9-r(1+\frac{\pi}{2})\).

OpenStudy (zehanz):

Why all this trouble? Well you have to minimize the area of the cross-section. This area is the area of a rectangle with dimensions 2r and h and a semicircle with radius r. Area of rectangle is 2r*h. Area of semicircle is pi*r²/2. Total area: \(A(r, h)=2rh+\frac{\pi}{2}r^2\). But: A is now a function of two variables. Because we now know that h depends on r, we don't need h. So, let's substitute \(h=9-r(1+\frac{\pi}{2})\) in the formula for A: \(A(r)=2r(9-r(1+\frac{\pi}{2}))+\frac{\pi}{2}r^2\). You have to simplify this formula as much as possible. When that is done, you can calculate A'(r) and solve A'(r)=0 to find the optimum.

OpenStudy (anonymous):

wait so the last equation is the original one, and when i calculate its derivative, it will give me maximum area of the cross section?

OpenStudy (zehanz):

Yes, after simplifying it, A'(r) = 0 gives the value of r with the largest cross-section.

OpenStudy (anonymous):

what is the derivative suppose to give me?

OpenStudy (zehanz):

\(A(r)=18r-2(1+\frac{\pi}{2})r^2+\frac{\pi}{2}r^2=18r-2r^2-\pi r^2+\frac{\pi}{2}r^2\). \(A(r)=18r-2r^2-\frac{\pi}{2}r^2=18r-(2+\frac{\pi}{2})r^2\). So it is a 2nd degree function. Then \(A'(r)=18-2(2+\frac{\pi}{2})r=0\) And: \(9=(2+\frac{\pi}{2})r\), So: \(r=\dfrac{9}{2+\frac{\pi}{2}}=\dfrac{9}{\frac{4+\pi}{2}}=9 \cdot \dfrac{2}{4+\pi}\). Now we found: \(r=\dfrac{18}{4+\pi}\). This will be the value of r that gives the maximum cross-section, although we could make a sign-scheme of A' to see that this r gives a maximum indeed.

OpenStudy (zehanz):

As you can see, such a short question generates a lot of work!

OpenStudy (anonymous):

yeah its tons of work! wait so the derivative of the equation is 18r−(2+π2)r2??

OpenStudy (zehanz):

Yes, if you mean π/2 by "π2".

OpenStudy (anonymous):

yeah sorry thats what i meant!

OpenStudy (anonymous):

and if i have to do the chart that has the value of -infinity and +infinity? like what is the range of the equation?

OpenStudy (anonymous):

wait were did you get the 9= part?

OpenStudy (zehanz):

By dividing by 2.

OpenStudy (anonymous):

ouu okok! got it! wow thank you so much!

OpenStudy (anonymous):

can you help me in another problem?

OpenStudy (zehanz):

Is your "Best Response" button stuck? :D

OpenStudy (zehanz):

Just kidding, I will help you anyway!

OpenStudy (anonymous):

ahahaha! ummm the problem is: Find the point on the graph of y=square root of x which is closest to the point (4,0)

OpenStudy (zehanz):

So this is the situation: (see image) You have to find the position of P, when AP is shortest.

OpenStudy (anonymous):

okk and how do i do that?

OpenStudy (zehanz):

You have to make a function that describes the length of AP. If P=(a, √a) and A=(4, 0), how do you calculate the length of AP?

OpenStudy (anonymous):

AP is the line right?

OpenStudy (anonymous):

i think i could do it with a triangle but I'm not sure

OpenStudy (zehanz):

See the image, AP is indeed the line you need. And yes, a triangle (and the Pythagorean Theorem) gives you the distance. Can you set up a formula?

OpenStudy (anonymous):

i know it has to have 4^2 but i don't really know how to get the other numbers

OpenStudy (anonymous):